OpenStudy (anonymous):

At a local school, 50% of the students take an advanced class, while 10% of the students participate in sports. What is the probability that a student selected at random takes an advanced class and participates in sports?

OpenStudy (anonymous):

@Kitt020912

OpenStudy (anonymous):

Kitt?

OpenStudy (anonymous):

Just a sec

OpenStudy (anonymous):

Okay:)

OpenStudy (anonymous):

\[50 percent \rightarrow \frac{ 50 }{ 100 }\rightarrow \frac{ 1 }{ 2 }\] \[10 percent \rightarrow \frac{ 10 }{ 100 }\rightarrow \frac{ 1 }{ 10}\] \[\frac{ 1}{ 2 }\times \frac{ 1}{ 10}=\frac{ 1 }{ 20 }\]

OpenStudy (anonymous):

TNX U!:0

OpenStudy (anonymous):

Im grateful. Thanks:)

OpenStudy (anonymous):

kitt, another question, can i tag u>?